OpenStudy (joehoffy):

Which of the following is a step in simplifying the expression {x*y^2/x^-3*x^3}^2

OpenStudy (joehoffy):

@Lovelarap

OpenStudy (anonymous):

Do you know that you have to multiply -4 to the rest of the powers?

OpenStudy (joehoffy):

?

OpenStudy (anonymous):

OK, let me draw it.

OpenStudy (anonymous):



OpenStudy (joehoffy):

lol what?

OpenStudy (anonymous):

So on the top, you always have power and square root. Then you have multiplication and division. Lastly plus and minus. When you have an equation, you always do power and square root first. Multiplication and division second. Plus and minus last.

OpenStudy (joehoffy):

okay can you walk me throu you doing it? becuase this is confusing

OpenStudy (anonymous):

Ok, basically you multiply the power. So multiply -4 to the rest of the powers in the bracket.

OpenStudy (joehoffy):

HOW? its xy^2!?

OpenStudy (anonymous):

The x is actually on the power of 1 and the y is on the power of 2. xy^2 is the same as x*y^2

OpenStudy (anonymous):

OK, let's do this: \[(\frac{ x^1 \times y^2 }{ x^-3 \times y^3 })^-4\]

OpenStudy (joehoffy):

how

OpenStudy (anonymous):

when you see xy or fg it means they are multiplying. So it is the same as x*y or f*g

OpenStudy (joehoffy):

ok but what do they stand for?